# Entanglement-breaking channels are a quantum memory resource

### v1

@@ Abstract
Entanglement-breaking channels (equivalently, measure-and-prepare channels) are an important class of quantum operations noted for their ability to destroy multipartite spatial quantum correlations. Inspired by this property, they have also been employed in defining notions of “classical memory”, under the assumption that such channels effectively act as a classical resource. We show that, in a single-system multi-time scenario, entanglement-breaking channels are still a quantum memory resource: a qudit going through an entanglement-breaking channel cannot be simulated by a classical system of same dimension. We provide explicit examples of memory-based output generation tasks where entanglement-breaking channels outperform classical memories of the same size. Our results imply that entanglement-breaking channels cannot be generally employed to characterize classical memory effects in temporal scenarios without additional assumptions.
Introduction.— The notion of classical memory, in the sense of storage of information, is familiar to us from our everyday life. When we talk about the number of bits (or, more commonly, one of its multiples: bytes, kilobytes, etc.) a device has, we refer the logarithm of the number of distinguishable internal states, where information can be stored and retrieved. Needless to say, memory is an essential resource for information processing, and a finite amount of memory imposes fundamental limitations on which tasks can be achieved. With the advent of quantum information and quantum computation, distinguishing between classical and quantum memory effects is of utmost importance for identifying scenarios where quantum systems offer an advantage over classical ones [1, 2, 3, 4]. An approach towards this goal is by establishing bounds on information processing tasks achievable with classical memory alone. A violation of such bounds then provides a certificate of non-classicality, often also providing a quantifiable metric for the quantumness of the underlying memory resource [5, 6, 7, 8, 9, 10, 11, 12].
Any information processing task consists of manipulating information stored in a memory over time. This is achieved by sequential operations performed on the memory, with classical and quantum systems offering vastly different capabilities on what memory states and their transformations are possible. In recent years, the formalism of quantum supermaps [13, 14, 15], also known as process tensors or quantum combs, has enabled a thorough investigation of temporal quantum protocols, as it is naturally well-suited for the study of multi-time statistics [7, 16, 17]. Within this framework, many quantum-inspired notions have been employed as conflicting definitions of “classical memory” [1, 16, 17, 4, 18]; see Ref. [17] for a detailed characterization.
@@
A typical approach involves entanglement-breaking (EB) channels, which are known to be exactly equivalent to measure-and-prepare (M&P) operations [19]. As the input quantum state is destroyed by the measurement operation and replaced by a classical label (albeit encoded in the prepared output quantum state), one may be inclined to claim that such channels are essentially equivalent to classical channels [1, 9, 10, 11, 16], i.e., that they can only establish classical correlations between their inputs and outputs. EB channels have been used in this manner, e.g., to develop resource theories of quantum memories [1], the benchmarking and application of quantum devices [9], characterizing non-classical memory effects in quantum processes [4, 20, 16] and bosonic Gaussian channels [21, 11], and in establishing a hierarchy between competing notions of classical memory [17]. A potential issue with such entanglement-inspired definitions is that they still operate within the quantum formalism, and therefore, may not truly characterize the limitations inherent to classical systems. Clearly, a precise distinction between classical and quantum memory effects must emerge from explicit constructions within each theory.
@@
Here, we show that the previous interpretation and applications of EB channels can be misleading in the case of temporal scenarios involving sequential measurements. We address the following question: is the memory of a “$d$-it” (a classical system with $d$ states) equivalent to that of a qudit ($d$-level quantum system) passing through an EB channel? We answer it in the negative by presenting explicit examples where a qudit passing through an EB channel outperforms a $d$-it of classical memory in the task of generating correlations in time. We conclude that EB channels cannot be used to define or characterize classical memories in the most general settings.
@@
Preliminary notions.— A channel $\mathcal{E}_{A}$ is said to be entanglement breaking if $(\mathcal{E}_{A}\otimes\text{id}_{B})(\rho_{AB})$ is separable for any bipartite state $\rho_{AB}$. A well-known result establishes that any EB channel $\mathcal{E}$ can be written as a measure-and-prepare (M&P) channel in the Holevo form [22, 19]
$\mathcal{E}(\rho)=\sum_{i=0}^{m}\operatorname{tr}\left[\rho E_{i}\right]\sigma_{i},$ (1)
@@
for $\{E_{i}\}_{i=1}^{m}$ a Positive Operator-Valued Measure (POVM), i.e. $E_{i}\geq 0$ and $\sum_{i=1}^{m}E_{i}=\mathbbm{1}$, and $\sigma_{i}$ quantum states. Here, $m$ can be interpreted as the number of labels—in principle, arbitrarily large—required in the channel’s intermediate classical memory in order to describe its operation [19, 1]. EB channels can be fully characterized [19] by the fact their Choi matrix [23, 24], $C_{\mathcal{E}}:=(\mathcal{E}\otimes\text{id})(\left|\Phi\right\rangle\left\langle\Phi\right|)$ with $\left|\Phi\right\rangle$ a maximally entangled state, is separable and with rank $r\geq d$, which also implies any such channel can be rewritten with at most $m=d^{2}$ operators $E_{i}$ and $\sigma_{i}$. These channels can be understood in terms of a quantum-classical-quantum operation: (1) The input quantum state $\rho$ is first measured, producing a classical outcome label $i$, out of a total $m$, with probability $\operatorname{tr}\left[\rho E_{i}\right]$, (2) Conditioned on $i$, the corresponding quantum state $\sigma_{i}$ is passed forward as output; see Fig. 2b.
@@
To properly distinguish between classical and quantum memories, we investigate their capabilities when generating temporal correlations in the simplest scenario, i.e., that of a single, finite-dimensional system subject to the same dichotomous measurement at multiple times, as shown in Fig. 1. Such a system can be described within the formalism of finite-state machines [6, 25]. In the classical case, the most general description of its behavior is by an initial stochastic vector $\pi$ over its $d$ distinct states (i.e., a $d$-it of memory), and a pair of (row) sub-stochastic transition matrices $T=(T_{0},T_{1})$ describing its dynamics upon each outcome $a\in\{0,1\}$, with $T_{0}+T_{1}$ row stochastic. The probability of a length-$L$ sequence of outcomes $\bm{a}=a_{1}a_{2}\cdots a_{L}$ is then given by
$p(\bm{a}|T,d)=\pi T_{a_{1}}\cdots T_{a_{L}}\eta,$ (2)
@@
with $\eta=(1,\dots,1)^{\text{T}}$. The quantum case is defined analogously, with a single $d$-dimensional quantum system (a qudit), initially at state $\rho_{0}$, being repeatedly measured by a quantum instrument $\mathcal{I}=(\mathcal{I}_{0},\mathcal{I}_{1})$, with effects $\mathcal{I}_{a}$ being completely positive (CP) trace non-increasing maps such that $\mathcal{I}_{0}+\mathcal{I}_{1}$ is completely positive trace-preserving (CPTP). The probability of a sequence $\bm{a}$ is given by
@@
In both cases, the system’s state is the sole resource available to generate the probability distributions, and any dynamics is entirely governed by the effects associated with an outcome $a_{t}$. In this way, classical and quantum memories are placed on equal footing as a memory resource. Crucially, memory is the central resource for this task, as any arbitrary temporal correlation can be generated by a classical system with enough memory, i.e., enough internal states [5, 26, 27]. Consequently, the system’s dimension establishes fundamental limitations on the set of achievable temporal correlations [28, 6, 27, 8, 25], providing a clear notion for distinguishing between manifestly classical and quantum memories. Note that we do not allow for external storage of information (e.g., by conditioning on past outcomes, or invoking an external reference clock), so that the total amount of memory can be quantified by the number of states $d$ of the system.
One possible approach to distinguish between classical and quantum memory effects is by establishing upper bounds on the maximum probability for individual sequences to be produced within classical and quantum mechanics [6, 8, 25]. The notion of deterministic complexity of a sequence $\bm{a}$, denoted by $\operatorname{DC}(\bm{a})$, was introduced in Ref. [25] as the minimum dimension $d$ for a system (classical or quantum) to be capable of generating $\bm{a}$ with probability one. Thus, if $d<\operatorname{DC}(\bm{a})$, then no classical or quantum system can produce the sequence $\bm{a}$ deterministically, implying both classical and quantum theory must obey non-trivial upper bounds for this maximum probability. Since $d=2$ is the smallest non-trivial amount of memory, these observations also highlight the importance of studying temporal correlations over three or more time steps, as every sequence of length two can be produced deterministically with either a bit or qubit.
As the underlying memory resource can be explicitly confined within each theory, such non-trivial upper bounds provide a sharp demarcation between classical and quantum memories and their capabilities, but they are generally difficult to obtain precisely [25, 29, 30]. In Ref. [25], a universal upper bound of $1/\mathrm{e}$ was conjectured for all sub-deterministic classical memory scenarios, whereas no non-trivial universal quantum bound seems to exist.
@@
The sequential measurement protocol.— We investigate the non-classicality of entanglement-breaking channels by inserting an EB channel $\mathcal{E}$, as in Eq. 1, before each quantum measurement in Eq. 3 (depicted in Fig. 2a), obtaining
@@
$\begin{split}&p(\bm{a}|\mathcal{I},\mathcal{E},d)=\operatorname{tr}\left[\mathcal{I}_{a_{L}}\circ\mathcal{E}\circ\cdots\circ\mathcal{I}_{a_{2}}\circ\mathcal{E}\circ\mathcal{I}_{a_{1}}\circ\mathcal{E}(\rho_{0})\right]\\ &=\sum_{i_{1}\dots i_{L}=1}^{m}\operatorname{tr}\left[\rho_{0}E_{i_{1}}\right]\operatorname{tr}\left[\mathcal{I}_{a_{1}}(\sigma_{i_{1}})E_{i_{2}}\right]\cdots\operatorname{tr}\left[\mathcal{I}_{a_{L}}(\sigma_{i_{L}})\mathbbm{1}\right]\\ &=\sum_{i_{1}\dots i_{L+1}=1}^{m}\pi_{i_{1}}[T_{a_{1}}]_{i_{1}i_{2}}\cdots[T_{a_{L}}]_{i_{L}i_{L+1}}\eta_{i_{L+1}},\end{split}$ (4)
@@
where we use $[T_{a}]_{ij}:=\operatorname{tr}\left[\mathcal{I}_{a}(\sigma_{i})E_{j}\right]$, $\pi_{i}:=\operatorname{tr}\left[\rho_{0}E_{i}\right]$, and $\eta=(1,\dots,1)^{\text{T}}$ as in Eq. 2. Note that, since the $E_{i}$ are POVM elements, each trace term denoted a probability. We have thus collected all such probabilities in the matrices $T_{a}$ such that
@@
$\sum_{a,j}[T_{a}]_{ij}=\operatorname{tr}\left[\textstyle{\sum_{a}}\mathcal{I}_{a}(\sigma_{i})\cdot\textstyle{\sum_{j}}E_{j}\right]=\operatorname{tr}\left[\hat{\sigma}_{i}\mathbbm{1}\right]=1,$ (5)
@@
where we use the fact $E_{i}$ form a POVM and $\sum_{a}\mathcal{I}_{a}$ is CPTP, thus $\hat{\sigma}_{i}=\sum_{a}\mathcal{I}_{a}(\sigma_{i})$ has unit trace. Furthermore, we can define an initial probability vector $\pi$ arising from the first application of the channel, with $\sum_{i}\pi_{i}=1$. The above results imply we can model the protocol classically, by recognizing $T_{a}$ as $m\times m$ sub-stochastic transition matrices and $\pi$ as an initial stochastic vector (Fig. 2c). Eq. 4 can thus be equivalently written as in Eq. 2, with $m$ instead of $d$ states.
@@
We are led to conclude that the repeated inclusion of an arbitrary entanglement-breaking channel before the quantum instrument allows the resulting temporal correlations to be, indeed, described classically. This observation seems to justify the association of EB channels with classical memories. Nevertheless, as we already discussed, any temporal correlation can be obtained via a classical system if enough memory is available, such that a meaningful comparison must be made between systems of bounded memory, i.e., bounded dimension.
@@
A central observation is that the effective classical model thus obtained is over $m$ “virtual” classical states (i.e., the number of outcomes in the measure-and-prepare description of the channel), not $d$ as in the dimension of the underlying Hilbert space. This suggests that a qudit passing through an EB channel may be able to generate stronger correlations, i.e., carry more memory, than a classical $d$-it. Clearly, the exception is the case $m\leq d$, where any quantum model can be converted into a classical one of the same dimension (and vice versa) by choosing $E_{i},\sigma_{i}$, and $\mathcal{I}$ acting diagonally on the same basis; see App. A for the explicit construction.
@@
Figure 3: (Color online). Probabilities of one-tick sequences $\bm{a}^{L}_{\text{ot}}$ for classical one-way models and the quantum construction using ETFs, both using $d=L-1$. One-way models are conjectured to be optimal for all $L=d-1$, and are known to achieve the exact upper bounds $\Omega_{C}(\bm{a}^{L}_{\text{ot}},d)$ for $L=3,4$, and $5$. The bounds for $L=4$ and $5$ are violated by this explicit quantum construction using ETFs, revealing non-classical memory effects of EB channels.
@@
Table 1: Results from numerical optimization for various one-tick sequences of length $L$, and Hilbert space dimension $d=L-1$. The quantum models found numerically for $L=4$ and $5$ outperform the previous ETF construction, and are included in App. D. No violations were found for $d=2$.
Quantum advantages with EB channels.— Here, we show that EB channels with $m>d$ are able to generate non-classical memory effects. This can be investigated by choosing a sequence $\bm{a}$ with $d<\operatorname{DC}(\bm{a})$, then computing the maximum probability achievable over all classical models with $d$ states:

### v2

@@ Abstract
Entanglement-breaking channels (equivalently, measure-and-prepare channels) are quantum operations noted for their ability to destroy multipartite spatial quantum correlations. Inspired by this property, they have also been widely employed in defining notions of “classical memory”, under the assumption that such channels effectively act as a classical resource. In this paper, we show this assumption is false. By means of multi-time correlations, we conclusively show that entanglement-breaking channels are still a quantum resource: a qudit going through an entanglement-breaking channel can generate genuinely nonclassical temporal correlations, i.e., that cannot be simulated by a classical system of same dimension. Our results imply that entanglement-breaking channels cannot generally be employed to characterize classical memory effects without additional assumptions.
I Introduction
Classical memory, in the sense of storage of information, is a familiar notion to us from daily life. With the advent of quantum information and quantum computation, several analogous notions of quantum memory have now been introduced to describe quantum resources offering advantages over their classical counterparts. These advantages have been attributed to disparate quantum phenomena in a variety of scenarios [1, 2, 3, 4, 5, 6, 7], making the essential features for a quantum memory the subject of ongoing debate. Nevertheless, even if we lack definitive criteria for the “quantumness” of a memory, its “classicality” should unambiguously indicate a total absence of any quantum advantages. Thus, a complete understanding of classical memory effects becomes essential to any practical notion of nonclassical memory effects resulting from a quantum memory.
Quantum memories feature prominently in the study of quantum networks [8, 9, 10, 11] and multi-time quantum processes [12, 13, 14]. The notion of classical and quantum memory also plays a role in open quantum systems [15, 16], where the environment acts as the memory resource leading to temporally-correlated noise, and in notions of classical and quantum communication capacities [17]. Clearly, a precise characterization of classical and quantum memories is of great theoretical and practical interest. To this end, the formalism of quantum supermaps [18, 19, 20] (also known as process tensors or quantum combs) has been extensively applied to the investigation of quantum processes [12, 13, 14, 16], where arbitrary protocols can be reinterpreted as multipartite quantum states. For this reason, several notions of “classical memory” have been proposed in analogy to spatial correlations, often in conflict with one another [14], but nevertheless all based on interpreting separability of a supermap as a signature of the classicality of its memory [8, 13, 11, 14, 21].
@@ I Introduction
Building on this, the typical approach to render the memory in a quantum supermap “classical” employs entanglement-breaking (EB) channels, known to be exactly equivalent to measure-and-prepare operations [22]. As the input quantum state is destroyed by the measurement and replaced by a classical label (albeit encoded in the prepared output quantum state), one may be inclined to claim that such channels are essentially equivalent to classical channels [8, 9, 23, 24, 25, 13, 26], i.e., that they can only transmit classical information between their inputs and outputs, therefore being restricted to classical correlations. EB channels have been used in this manner, e.g., to develop resource theories of quantum memories [8, 13], the benchmarking of quantum devices [23], characterizing memory effects in quantum processes [11, 27, 13, 28, 26] and bosonic Gaussian channels [29, 25], and in establishing a hierarchy between competing notions of quantum processes with classical memory [14].
The issue with these approaches is that they still operate within the quantum formalism, obscuring the distinction between quantum and classical memory effects, and thereby preventing a proper characterization of these memory resources. Instead, the distinction between classical and quantum memory effects should emerge from explicit constructions within each theory. Thus, a more rigorous approach consists in establishing bounds on information processing tasks achievable with classical memory alone. A violation of these bounds then provides a certificate of nonclassicality, often also providing a quantifiable metric for the quantumness of the underlying memory resource. Crucially, this requires not only the most general possible notion of “classical memory”, but also an appropriate choice of a task where classical and quantum memories can be compared on equal terms.
@@
In this paper, we show that interpreting EB channels generically as classical memory resources can be misguided. We address the following question: is the memory of a dit (a classical system with $d$ states) equivalent to that of a qudit ($d$-level quantum system) passing through an EB channel? We answer it in the negative, by presenting explicit scenarios where a qudit passing through an EB channel outperforms a dit of classical memory in the task of generating correlations in time. We conclude that, contrary to prevailing interpretations and applications, EB channels cannot generically be used to define or characterize classical memory effects in multi-time scenarios.
@@ II Preliminary notions
A channel $\mathcal{E}_{A}$ is said to be entanglement-breaking if $(\mathcal{E}_{A}\otimes\text{id}_{B})(\rho_{AB})$ is separable for any bipartite state $\rho_{AB}$. A well-known result establishes that any EB channel $\mathcal{E}$ can be written as a measure-and-prepare channel in the Holevo form [30, 22]
$\mathcal{E}(\rho)=\sum_{i=1}^{m}\operatorname{tr}\left[\rho E_{i}\right]\sigma_{i},$ (1)
@@
for $\{E_{i}\}_{i=1}^{m}$ a positive operator-valued measure (POVM), i.e., $E_{i}\geq 0$ and $\sum_{i=1}^{m}E_{i}=\mathbbm{1}$, and $\sigma_{i}$ quantum states. Here, $m$ can be interpreted as the number of labels—in principle, arbitrarily large—required in the channel’s intermediate classical memory in order to describe its operation [22, 8]. These channels can be understood in terms of a quantum-classical-quantum operation: (1) The input quantum state $\rho$ is first measured, producing a classical outcome label $i$, out of a total $m$, with probability $\operatorname{tr}\left[\rho E_{i}\right]$, (2) Conditioned on $i$, the corresponding quantum state $\sigma_{i}$ is passed forward as output; see Fig. 2b. EB channels can be fully characterized [22] by the fact their Choi matrix [31, 32], $C_{\mathcal{E}}:=(\mathcal{E}\otimes\text{id})(\left|\Phi\right\rangle\left\langle\Phi\right|)$ with $\left|\Phi\right\rangle$ a maximally entangled state, is separable and with rank $r\geq d$, which also implies any such channel can be rewritten with at most $m=d^{2}$ operators $E_{i}$ and $\sigma_{i}$.
@@
To precisely distinguish between classical and quantum memories, we investigate their capabilities when generating temporal correlations in the simplest scenario, i.e., that of a single, finite-dimensional system subject to the same dichotomous measurement at multiple times, as shown in Fig.˜1. This scenario can be described within the formalism of finite-state machines [33, 34, 35], recently introduced to the study of temporal correlations as it allows an exact characterization of classical, quantum, and generalized probability theories [33].
In the classical case, the most general finite-state machine description is given by an initial stochastic vector $\pi$ over its $d$ distinct states (i.e., a dit of memory), and a pair of (column) sub-stochastic transition matrices $T=(T_{0},T_{1})$ describing its dynamics upon each outcome $a\in\{0,1\}$, with $T_{0}+T_{1}$ column stochastic. Importantly, this description allows classical measurements to be invasive, going beyond typical definitions of classicality rooted on unnecessarily macroscopic notions of realism, as in Leggett-Garg inequalities [36, 37, 38, 39] and in the application of the Kolmogorov extension theorem to stochastic processes [40]. The probability of a length-$L$ sequence of outcomes $\bm{a}=a_{1}a_{2}\ldots a_{L}$ is then given by
$p(\bm{a}|T,d)=\eta T_{a_{L}}\cdots T_{a_{2}}T_{a_{1}}\pi,$ (2)
@@
with $\eta=(1,\dots,1)$. The quantum case is defined analogously, with a single $d$-dimensional quantum system (a qudit), initially at state $\rho_{0}$, being repeatedly measured by a quantum instrument $\mathcal{I}=(\mathcal{I}_{0},\mathcal{I}_{1})$, with effects $\mathcal{I}_{a}$ being completely positive (CP) trace non-increasing maps such that $\mathcal{I}_{0}+\mathcal{I}_{1}$ is completely positive trace-preserving (CPTP). The probability of a sequence $\bm{a}$ is given by
@@
In both classical and quantum cases, the system’s state is the sole memory resource available to generate the probability distributions, and any dynamics is entirely governed by the effects associated with a measurement outcome $a_{t}$. In this way, classical and quantum memories are placed on equal terms as a memory resource.
A key observation is that memory is the main resource for realizing the distributions in Eqs.˜2 and 3, as any temporal correlation can be generated by a classical or quantum system with enough memory, i.e., enough internal states [41, 42, 43, 44]. Consequently, the system’s dimension establishes fundamental limitations on the set of achievable temporal correlations [43, 33, 44, 34, 35], providing an unambiguous notion for distinguishing between classical and quantum memories. In other words, bounding the amount of memory is absolutely essential for establishing any meaningful notion of classicality or nonclassicality of temporal correlations. Note that we do not allow external storage of information (e.g., by conditioning on past outcomes, the use of shared randomness between time steps, or invoking an external reference clock), ensuring that the total amount of memory can be precisely accounted for by the number of orthogonal states $d$ of the memory system.
One way to distinguish between classical and quantum memory effects is by establishing upper bounds on the task of sequence generation, i.e., the maximum probability for individual sequences to be produced within classical and quantum theory [33, 34, 35]. The notion of deterministic complexity of a sequence $\bm{a}$, denoted by $\operatorname{DC}(\bm{a})$, was introduced in Ref. [35] as the minimum dimension $d$ for a system (classical or quantum) to be capable of generating $\bm{a}$ with probability one. Thus, if $d<\operatorname{DC}(\bm{a})$, then no classical or quantum system can produce the sequence $\bm{a}$ deterministically, implying both classical and quantum theory must obey nontrivial upper bounds for the maximum probability. Since $d=2$ is the smallest nontrivial amount of memory, these observations also highlight the importance of studying temporal correlations—and thus, memory effects—in scenarios with three or more time steps, as every sequence of length two can be produced deterministically with either a bit or qubit.
As the underlying memory resource can be explicitly confined within each theory, such nontrivial upper bounds provide a sharp demarcation between classical and quantum memories and their capabilities, but they are generally difficult to obtain precisely [35, 45, 46]. In Ref. [35], a universal upper bound of $1/\mathrm{e}$ was conjectured for all sub-deterministic classical memory scenarios, whereas no nontrivial universal quantum bound seems to exist.
@@ III The sequential measurement protocol
We investigate the nonclassicality of entanglement-breaking channels by inserting a fixed EB channel $\mathcal{E}$, as in Eq.˜1, before each quantum measurement in Eq.˜3 (depicted in Fig. 2a), obtaining
@@
$\begin{split}&p(\bm{a}|\mathcal{I},\mathcal{E},d)=\operatorname{tr}\left[\mathcal{I}_{a_{L}}\circ\mathcal{E}\circ\cdots\circ\mathcal{I}_{a_{2}}\circ\mathcal{E}\circ\mathcal{I}_{a_{1}}\circ\mathcal{E}(\rho_{0})\right]\\ &=\sum_{i_{1},\ldots,i_{L}=1}^{m}\operatorname{tr}\left[\mathcal{I}_{a_{L}}(\sigma_{i_{L}})\mathbbm{1}\right]\cdots\operatorname{tr}\left[\mathcal{I}_{a_{1}}(\sigma_{i_{1}})E_{i_{2}}\right]\operatorname{tr}\left[\rho_{0}E_{i_{1}}\right]\\ &=\sum_{i_{1},\ldots,i_{L+1}=1}^{m}\eta_{i_{L+1}}[T_{a_{L}}]_{i_{L+1}i_{L}}\cdots[T_{a_{1}}]_{i_{2}i_{1}}\pi_{i_{1}},\end{split}$ (4)
@@
where we use $T_{a}]_{ji}:=\operatorname{tr}\left[\mathcal{I}_{a}(\sigma_{i})E_{j}\right]$, $\pi_{i}:=\operatorname{tr}\left[\rho_{0}E_{i}\right]$, and $\eta=(1,\dots,1)$ as in Eq.˜2. Note that, since the $E_{i}$ are POVM elements, each trace term corresponds to a transition probability. We have thus collected all such probabilities in the matrices $T_{a}$ such that
@@
$\sum_{a,j}[T_{a}]_{ji}=\operatorname{tr}\left[\textstyle{\sum_{a}}\mathcal{I}_{a}(\sigma_{i})\cdot\textstyle{\sum_{j}}E_{j}\right]=\operatorname{tr}\left[\hat{\sigma}_{i}\mathbbm{1}\right]=1,$ (5)
@@
where we use the fact $E_{i}$ form a POVM and $\sum_{a}\mathcal{I}_{a}$ is CPTP, thus $\hat{\sigma}_{i}=\sum_{a}\mathcal{I}_{a}(\sigma_{i})$ has unit trace. Furthermore, we can define an initial probability vector $\pi$ arising from the first application of the channel, with $\sum_{i}\pi_{i}=1$. The above results imply we can model the protocol classically, by recognizing $T_{a}$ as $m\times m$ sub-stochastic transition matrices and $\pi$ as an initial stochastic vector (Fig. 2c). Eq.˜4 can thus be equivalently written as in Eq.˜2, with $m$ instead of $d$ states.
@@
We are led to conclude that the repeated inclusion of an arbitrary entanglement-breaking channel before the quantum instrument allows the resulting temporal correlations to be, indeed, described classically. This observation seems to further justify the association of EB channels with classical memories. However, as we already discussed, any temporal correlation can be obtained via a classical system if enough memory is available, rendering the previous observation vacuous unless the underlying amount of memory is also taken into account. Thus, any meaningful comparison must be made between systems of bounded memory, i.e., bounded dimension.
@@
Here, a key observation is that the effective classical model obtained in Eq.˜4 is over $m$ classical “virtual” states (i.e., the number of outcomes in the measure-and-prepare description of the $\mathcal{E}$), not $d$ as in the dimension of the underlying quantum system. This suggests that a qudit passing through an EB channel may be able to generate stronger correlations than a classical dit. Clearly, the exception is the case $m\leq d$, where any quantum model can be converted into a classical one of the same dimension (and vice versa) by choosing $E_{i}$, $\sigma_{i}$, and $\mathcal{I}$ acting diagonally on the same basis; see Appendix˜A for the explicit construction.
@@
Figure 3: Probabilities of one-tick sequences $\bm{a}^{L}_{\text{ot}}$ for classical one-way models (cf. Ref. [34, 35]) and the quantum construction using ETFs, both using d=L1d=L-1. One-way models are conjectured to be optimal for all L=d1L=d-1, and are known to achieve the exact upper bounds $\Omega_{C}(\bm{a}^{L}_{\text{ot}},d)$ for $L=3,4$, and $5$. The bounds for $L=4$ and $5$ are violated by this explicit quantum construction using ETFs, revealing nonclassical memory effects of EB channels.
@@
Table 1: Results from numerical optimization for various one-tick sequences of length $L$, and Hilbert space dimension d=L1d=L-1. The quantum models found numerically for $L=4$ and 55 outperform the previous ETF construction, and are included in Appendix˜C. No violations were found for $d=2$.
IV Quantum advantages with entanglement-breaking channels
Here, we show that EB channels with $m>d$ are able to generate nonclassical memory effects. This can be investigated by choosing a sequence $\bm{a}$ with $d<\operatorname{DC}(\bm{a})$, then computing the maximum probability achievable over all classical models with $d$ states: